OpenStudy (anonymous):

If the equation y = -10 + 2.5x was plotted: Answer the vertical intercept would be -10. the slope would be -7.5. it would graph as a downsloping line. the slope would be -10.

OpenStudy (anonymous):

[y =-10+2.5x then the slope will be 2.5 and the y intercept is -10 x intercept is 4 since the slope is positive so it has alwyas positive slope

OpenStudy (anonymous):

Equation is of the form y = mx + b where m is the slope and b is the y (vertical) intercept. m = 2.5 slope b = -10 y (vertical) intercept dividing both sides by 2.5 yields y/2.5 = -4 + x when y=0 x=4 (x intercept)